# for Verifiable Chain-of-Thought Financial Reasoning

### v2

@@ Abstract
Multi-step symbolic reasoning is essential for robust financial analysis; yet, current benchmarks largely overlook this capability. Existing datasets such as FinQA and ConvFinQA emphasize final numerical answers while neglecting the intermediate reasoning required for transparency and verification. To address this gap, we introduce FinChain, the first benchmark specifically designed for verifiable Chain-of-Thought (CoT) evaluation in finance. FinChain spans 58 topics across 12 financial domains, each represented by parameterized symbolic templates with executable Python traces that enable fully machine-verifiable reasoning and scalable, contamination-free data generation. To assess reasoning capacity, we propose ChainEval, a dynamic alignment metric that jointly evaluates both the final-answer correctness and the step-level reasoning consistency. Evaluating 26 leading LLMs reveals that even frontier proprietary systems exhibit clear limitations in symbolic financial reasoning, while domain-adapted and math-enhanced fine-tuned models substantially narrow this gap. Overall, FinChain exposes persistent weaknesses in multi-step financial reasoning and provides a foundation for developing trustworthy, interpretable, and verifiable financial AI.
@@
Zhuohan Xie1 Daniil Orel1 Rushil Thareja1 Dhruv Sahnan1 Hachem Madmoun1,2 Fan Zhang3 Debopriyo Banerjee1 Georgi Georgiev4 Xueqing Peng5 Lingfei Qian5 Jimin Huang5 Jinyan Su6 Aaryamonvikram Singh1 Rui Xing7 Rania Elbadry1 Chen Xu5 Haonan Li1 Fajri Koto1 Ivan Koychev4 Tanmoy Chakraborty8 Yuxia Wang9 Salem Lahlou1 Veselin Stoyanov1 Sophia Ananiadou10 Preslav Nakov1 1MBZUAI 2Syllogia 3The University of Tokyo 4Sofia University “St. Kliment Ohridski” 5The Fin AI 6Cornell University 7The University of Melbourne 8IIT Delhi 9INSAIT 10The University of Manchester {zhuohan.xie, preslav.nakov}@mbzuai.ac.ae xueqing.peng2023@gmail.com ProjectCode  5 Leaderboard
**footnotetext: Co-second author.
@@ 1 Introduction
Large language models (LLMs) have shown strong performance across diverse tasks (Zhao et al., 2023a; Xie et al., 2023b), including applications in finance, healthcare, and law (Chen et al., 2024b). In finance, effective analysis often requires processing large volumes of text from reports, news, and social media, which frequently reflect or influence financial phenomena such as market sentiment (Nie et al., 2024).
FinChain Template (Compound Interest) #Question:
@@
Prior work in financial NLP has mainly adapted general-purpose tasks to financial documents, including information extraction (Shah et al., 2023), sentiment analysis (Pei et al., 2022), and text classification (Sy et al., 2023), where models produce relatively simple targets such as entity spans or sentiment labels. In contrast, financial reasoning requires generating chain-of-thought (CoT) traces that justify each step in solving a problem, as illustrated in Figure 1.
Existing benchmarks such as FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022) frame reasoning as numerical question answering, emphasizing final answers without requiring explicit or verifiable intermediate reasoning. Although some examples include partial traces, these are neither comprehensive nor rigorously structured. Lacking explicit step-level supervision limits diagnostic power: such benchmarks cannot reveal where reasoning fails or distinguish genuine multi-step inference from pattern matching. In mathematical reasoning, the symbolic-template paradigm introduced by GSM-Symbolic (Mirzadeh et al., 2024) re-templates problems from GSM8K (Cobbe et al., 2021) to provide explicit intermediate steps. We adopt a similar approach but construct our dataset entirely from scratch for finance. As shown in Figure 1, each symbolic template encodes a parameterized financial problem (e.g., compound interest) with named variables and numeric inputs, paired with executable Python code computing both intermediate steps and final results. This design enables scalable, contamination-free generation of consistent examples for training and evaluation. The financial domain spans diverse topics with distinct reasoning styles. To capture this diversity, we organize our dataset using a fine-grained taxonomy (Figure 2) covering 12 domains (e.g., Corporate Finance, Sustainable Finance, Crypto) and 58 topics. For each topic, we develop five parameterized templates: two easy, two intermediate, and one advanced, varying in symbolic complexity and required expertise. This forms the most detailed taxonomy of financial reasoning tasks to date. Each templated instance includes (1) a scenario card describing the topic, difficulty, and sampled inputs (e.g., “Discounted cash flow valuation, advanced”) and (2) an executable chain of reasoning steps in Python grounded in domain-specific formulas. Because every operation is explicit and executable, the benchmark is fully machine-verifiable: hallucinated, skipped, or incorrect steps can be automatically detected. This contrasts with datasets like FinQA and ConvFinQA, which supervise only final answers.
To enable rigorous and interpretable evaluation, we introduce ChainEval, a dynamic-alignment metric that jointly measures final-answer correctness and intermediate step faithfulness. Unlike conventional text similarity metrics, ChainEval explicitly verifies both semantic and numerical alignment between gold and predicted reasoning chains, providing a faithful reflection of symbolic reasoning quality. Our large scale evaluation across proprietary and open LLMs reveals a consistent pattern: frontier systems such as GPT-5, Claude 4.5, Gemini 2.5 Pro, and the subset evaluated Grok 4 Heavy achieve the highest overall accuracy, yet still exhibit systematic weaknesses in long horizon, compositional reasoning. Meanwhile, compact or domain specific models improve through fine-tuning but remain limited in multi-step symbolic inference, indicating that robust financial reasoning requires both domain grounding and sufficient capacity for structured, multi-hop abstraction. Our main contributions are as follows:
@@
  • We propose ChainEval, a verifiable reasoning metric that evaluates both step-level consistency and final-answer correctness, showing the strongest correlation with expert human judgments.
  • We benchmark 26 leading proprietary and open LLMs, including frontier, math-enhanced, and finance-tuned systems, revealing that even SOTA models struggle with verifiable multi-step reasoning, especially on advanced symbolic templates requiring compositional and numerical abstraction.
@@ 2.1 Financial NLP
Progress in financial NLP has been driven by both modeling and benchmarking. Early work focused on extraction and classification with models such as FinBERT (Araci, 2019), while later efforts expanded to personal finance (Hean et al., 2025), credit scoring (Feng et al., 2023), and risk-awareness benchmarking (Yuan et al., 2024). Datasets like FiNER-ORD, REFinD, FinARG, and ECTSum support tasks in NER, relation extraction, argument mining, and summarization (Shah et al., 2023; Kaur et al., 2023; Mukherjee et al., 2022; Xie et al., 2024). Large financial language models have further advanced the field. BloombergGPT (Wu et al., 2023) achieved broad domain performance, FinGPT (Liu et al., 2023) emphasized open-source adaptability, and FinMA (Xie et al., 2023a) delivered competitive results with a compact architecture. Corresponding benchmarks such as FLANG (Shah et al., 2022), FinBen (Xie et al., 2024), and FinMTEB (Tang and Yang, 2025) broadened evaluation coverage across diverse tasks, while BizBench (Koncel-Kedziorski et al., 2023) and PIXIU (Xie et al., 2023a) revealed limitations in quantitative and multimodal reasoning.
Despite this progress, limitations remain in multi-step reasoning, long-context understanding, and cross-market generalization (Chen et al., 2024b). These challenges motivate the need for benchmarks that assess capability of models to perform faithful, auditable reasoning grounded in financial knowledge.
@@ 2.2 Financial Reasoning
Real-world financial problems require precise numerical reasoning. FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022) supervise arithmetic program generation but provide only weak step-level supervision, yielding traces that are neither explicit nor verifiable. FinTextQA (Chen et al., 2024a) introduces long-form financial questions from textbooks and regulatory sources, focusing on explanatory retrieval rather than traceable computation. Bridging text and numerical reasoning, TAT-QA (Zhu et al., 2021) and MultiHiertt (Zhao et al., 2022) combine textual and tabular evidence, while DocMath-Eval (Zhao et al., 2023c) and FinanceMath (Zhao et al., 2023b) move toward interpretable, symbolic evaluation. However, these datasets still lack explicit, step-level, domain-aware supervision within authentic financial contexts. Model-centric advances such as FinR1 (Liu et al., 2025) and Fino1 (Qian et al., 2025) improve performance through program-based reinforcement learning and instruction tuning but remain limited in multi-step reasoning.
Overall, existing benchmarks offer either coarse program supervision (FinQA, ConvFinQA), fine-grained but domain-agnostic traces (GSM-Symbolic, DocMath-Eval), or retrieval-based QA without traceability (FinTextQA, TAT-QA, MultiHiertt, FinanceMath). None provide systematic, step-level supervision grounded in financial equations. FinChain fills this gap by introducing a symbolic, executable benchmark that supervises each intermediate step, supports automatic alignment and human assessment, and spans 58 topics across 12 financial domains.
@@ 3.1 Data Creation Process
We begin by identifying and defining financial domains based on established literature (Bodie et al., 2025) and expert input within the team. This process yields 12 distinct domains. To generate topics within each domain, we extract relevant passages from the literature and prompt ChatGPT  [^1] We use GPT-4o, the latest version available at the time of writing. with the domain name to propose candidate financial topics. Financial experts then curate these outputs, resulting in 58 topics in total (mean 4.8 per domain). The resulting taxonomy is illustrated in Figure 2. Following Mirzadeh et al. (2024), we implement executable Python methods that generate questions and solutions with chain-of-thought reasoning using ChatGPT. We upload the full paper and prompt ChatGPT to summarize the data creation process, then instruct it to produce ten template based questions per topic, explicitly varying complexity (easy, intermediate, and advanced) according to the number of reasoning steps. The template creation prompt is provided in Appendix A. Finally, we manually select five representative templates per topic (two easy, two intermediate, and one advanced) to ensure balanced difficulty in FinChain.
3.2 Data Validation
Directly prompting ChatGPT to mimic our style and generate questions sometimes fails. After iteratively inspecting generations and refining prompts, we introduced additional constraints to address recurring issues:
Cross-national inconsistencies.
Prompts occasionally produced hyper-localized financial contexts (e.g., currencies, exchange rates, indices, terminology from various countries). We standardized all questions to U.S.-based financial settings.
Precision mismatch.
Questions and solutions often displayed rounded values while computations used full precision. We aligned computational outputs with the displayed precision.
Incomplete input specification.
Some questions omitted variables required for calculation. We edited these cases to include all necessary inputs.
Unit consistency.
Currency notation (e.g., \$) was inconsistently applied across questions and solutions. We standardized units throughout.
Non-informative steps.
Certain solutions decomposed a single calculation into trivial substeps or only presented results. We revised such cases to reflect substantive reasoning steps.
Multiple targets.
Some questions requested multiple values, which complicates the evaluation process. We constrained questions to require a single target value.
3.3 Expert Check
To further enhance data quality, we recruited ten financial experts who volunteered to participate in this project. The team included seven graduate students in economics, finance, and related quantitative disciplines, and three industry professionals with practical experience in quantitative research, financial engineering, and risk management. All reviewers were carefully selected through an internal vetting process to ensure domain expertise and professional credibility. Their diverse profiles across academia, asset management, auditing, and financial technology helped maintain annotation consistency and domain coverage. Further demographic and professional details of the annotators are provided in Appendix B. These experts were tasked with carefully reviewing the ChatGPT-generated templates after the data validation phase. Before starting the main annotation, we conducted a pilot study to ensure consistent evaluation standards among annotators, followed by the distribution of the remaining tasks.
Annotation Platform
To facilitate efficient and user-friendly annotation, we developed a Streamlit-based annotation platform. Implementation details are presented in Appendix C.
Pilot Study
We randomly selected 20 templates for the pilot study and asked all annotators to review them. The goal was to align their annotation standards. After calibration, all annotators agreed that the selected templates were correct.
Main Annotation
After the pilot study, the remaining 270 templates were randomly distributed among the annotators, with each template reviewed by a single expert. Out of the total 290 templates, 29 were identified as incorrect. For those templates, we use the versions which are fixed by our financial experts. Detailed statistics of the identified issues are summarized in Appendix D.
@@ 4 ChainEval
We propose ChainEval, an evaluation framework assessing model outputs along two axes: final-answer correctness and reasoning-step alignment. Inspired by prior work on reasoning consistency (Lyu et al., 2023; Golovneva et al., 2023), we extend it by explicitly verifying intermediate results through step–answer matching. Unlike earlier approaches that focus mainly on textual consistency, our framework also checks the final numeric answer, providing a holistic view of model performance in both reasoning faithfulness and end-task fidelity. To our knowledge, this combination has not been explored in financial reasoning.
@@ 4.1 Preliminaries
where $s^{*}_{i}$ and $\hat{s}_{j}$ denote individual reasoning steps in the gold and predicted solutions. Each step $s_{i}$ produces an intermediate result,
@@
representing the numerical or symbolic value computed at that step.
To evaluate reasoning faithfulness, we compare these sequences both semantically and numerically. In addition, we employ Dynamic Time Warping (DTW) to capture the global structural alignment between step sequences. DTW provides an order-preserving but flexible alignment that accommodates insertions, deletions, or small reordering of steps while maintaining overall sequence coherence.
@@ 4.2 Reasoning Step Alignment
Each step is encoded using a sentence encoder $\mathrm{Enc}(\cdot)$, and pairwise semantic similarity between gold and predicted steps is computed as
$\mathrm{SS}(s^{*}_{i},\hat{s}_{j})=\mathrm{CosSim}\!\left(\mathrm{Enc}(s^{*}_{i}),\mathrm{Enc}(\hat{s}_{j})\right),$ (3)
where $\mathrm{CosSim}(\cdot,\cdot)$ denotes cosine similarity and $\mathrm{SS}\in[0,1]$.
@@
where $\mathbb{I}(\cdot)$ is the indicator function, and $\epsilon=0.05$ allows for up to 5% relative numeric deviation due to rounding or propagation errors.
@@
To capture global reasoning consistency, we align the two step sequences using Dynamic Time Warping (DTW). DTW searches for an optimal monotonic alignment path between (S,S^)($^{*},\hat{S}$) that minimizes cumulative cost while respecting step order. This formulation naturally handles local insertions, deletions, and step compressions, providing a structured measure of how closely the reasoning flows align.
Normalized DTW Alignment Score (Gate Mode).
We transform the minimal DTW cost into a normalized similarity measure:
$\mathrm{DTWNormGate}(S^{*},\hat{S})=1-\frac{\mathrm{Cost}_{\mathrm{DTW}}}{L_{\mathrm{path}}},$ (6)
where $\mathrm{Cost}_{\mathrm{DTW}}$ denotes the total alignment cost and $L_{\mathrm{path}}$ represents the length of the optimal alignment path. The resulting score lies in the range $[0,1]$, with higher values indicating stronger reasoning alignment between the gold and predicted solutions.

### v3

@@ Abstract
Multi-step symbolic reasoning is essential for robust financial analysis; yet, current benchmarks largely overlook this capability. Existing datasets such as FinQA and ConvFinQA emphasize final numerical answers while neglecting the intermediate reasoning required for transparency and verification. To address this gap, we introduce FinChain, the first benchmark specifically designed for verifiable Chain-of-Thought (CoT) evaluation in finance. FinChain spans 58 topics across 12 financial domains, each represented by parameterized symbolic templates with executable Python traces that enable fully machine-verifiable reasoning and scalable, contamination-free data generation. To assess reasoning capacity, we propose ChainEval, a dynamic alignment measure that jointly evaluates both the final-answer correctness and the step-level reasoning consistency. Our evaluation of 26 leading LLMs reveals that even frontier proprietary LLMs exhibit clear limitations in symbolic financial reasoning, while domain-adapted and math-enhanced fine-tuned models can substantially narrow this gap. Overall, FinChain exposes persistent weaknesses in multi-step financial reasoning and provides a foundation for developing trustworthy, interpretable, and verifiable financial AI.
@@
Zhuohan Xie1 Daniil Orel1thanks: Co-second author Rushil Thareja [^1] footnotemark: 1 Dhruv Sahnan [^1] footnotemark: 1 Hachem Madmoun1,2 Fan Zhang3 Debopriyo Banerjee1 Georgi Georgiev4 Xueqing Peng5thanks: Corresponding author Lingfei Qian5 Jimin Huang5 Jinyan Su6 Aaryamonvikram Singh1 Rui Xing7 Rania Elbadry1 Chen Xu5 Haonan Li1 Fajri Koto1 Ivan Koychev4 Tanmoy Chakraborty8 Yuxia Wang9 Salem Lahlou1 Veselin Stoyanov1 Sophia Ananiadou10 Preslav Nakov1 1MBZUAI 2Syllogia 3The University of Tokyo 4Sofia University “St. Kliment Ohridski” 5The Fin AI 6Cornell University 7The University of Melbourne 8IIT Delhi 9INSAIT 10The University of Manchester {zhuohan.xie, preslav.nakov}@mbzuai.ac.ae xueqing.peng2023@gmail.com Project Code Leaderboard
@@ 1 Introduction
Large language models (LLMs) have demonstrated strong performance across a wide range of tasks (Zhao et al., 2023a; Xie et al., 2023b), including applications in finance (Chen et al., 2024b). In this domain, effective analysis often requires synthesizing large volumes of textual information from reports, news, and social media, which reflect and influence financial phenomena such as market sentiment (Nie et al., 2024).
FinChain (Compound Interest) #Question:
@@
Most prior work in financial NLP has focused on tasks with shallow supervision, including information extraction (Shah et al., 2023), sentiment analysis (Pei et al., 2022), and text classification (Sy et al., 2023). These tasks typically require models to produce short outputs and do not test whether they can perform transparent, multi-step financial reasoning. In contrast, many financial problems require generating structured chains of reasoning that justify each intermediate step, as illustrated in Figure 1.
Existing financial reasoning benchmarks such as FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022) primarily frame reasoning as a numerical question answering and emphasize final-answer correctness. While some examples include intermediate traces, these are neither systematically structured nor rigorously verified. As a result, such benchmarks cannot reliably diagnose where reasoning fails or distinguish genuine multi-step inference from surface-level pattern matching.
Inspired by the symbolic-template paradigm introduced in mathematical reasoning (Mirzadeh et al., 2024), we construct a financial reasoning benchmark entirely from scratch. As shown in Figure 1, each symbolic template encodes a parameterized financial problem with named variables and numeric inputs, paired with executable Python code that computes both intermediate steps and final results. This design supports scalable, contamination-free data generation grounded in explicit symbolic and numerical operations. Financial reasoning spans diverse topics and requires heterogeneous expertise. To capture this diversity, we organize our dataset using a fine-grained taxonomy (see Figure 2) covering 12 domains and 58 topics. For each topic, we design five parameterized templates of increasing difficulty, comprising two easy, two intermediate, and one advanced template. Each instantiated example consists of a scenario card specifying the topic, the difficulty, and some example inputs, together with an executable chain of reasoning steps grounded in domain-specific formulas. Because the gold reasoning traces are explicit and executable, the intermediate computations and the final results can be verified at the symbolic and the numerical levels, thus enabling automatic detection of incorrect or inconsistent reasoning steps. To support rigorous and interpretable evaluation, we introduce ChainEval, a dynamic-alignment metric that jointly evaluates final-answer correctness and intermediate step faithfulness. Unlike conventional text similarity measures, ChainEval explicitly accounts for both semantic correspondence and numerical consistency between predicted and reference reasoning chains. Using this benchmark and evaluation framework, we evaluate 26 proprietary and open-weight LLMs. We find that frontier LLMs perform best overall, yet consistently struggle with advanced multi-step symbolic financial reasoning, while fine-tuned compact models achieve only limited gains.
Our main contributions are as follows:
@@
  • We propose ChainEval, a verifiable reasoning measure that evaluates both step-level consistency and final-answer correctness, and shows the strongest correlation with expert human judgments.
  • We benchmark 26 leading proprietary and open-weight LLMs, and find that even state-of-the-art LLMs struggle with verifiable multi-step financial reasoning, particularly on advanced symbolic templates.
@@ 2.1 Financial NLP
Progress in financial NLP has been driven by both modeling and benchmarking. Early work focused on extraction and classification with models such as FinBERT (Araci, 2019), while later efforts expanded to personal finance (Hean et al., 2025), credit scoring (Feng et al., 2023), and risk-awareness benchmarking (Yuan et al., 2024). Datasets like FiNER-ORD, REFinD, FinARG, and ECTSum support tasks in NER, relation extraction, argument mining, and summarization (Shah et al., 2023; Kaur et al., 2023; Mukherjee et al., 2022; Xie et al., 2024). Large financial language models have further advanced the field. BloombergGPT (Wu et al., 2023) achieved broad in-domain performance, FinGPT (Liu et al., 2023) emphasized open-source adaptability, and FinMA (Xie et al., 2023a) delivered competitive results with a compact architecture. Corresponding benchmarks such as FLANG (Shah et al., 2022), FinBen (Xie et al., 2024), and FinMTEB (Tang and Yang, 2025) broadened evaluation coverage across diverse tasks, while BizBench (Koncel-Kedziorski et al., 2023) and PIXIU (Xie et al., 2023a) revealed the limitations of quantitative and multimodal reasoning.
Despite this progress, limitations remain in multi-step reasoning, long-context understanding, and cross-market generalization (Chen et al., 2024b). These challenges motivate the need for benchmarks that assess the capability of LLMs to perform faithful, auditable reasoning that is grounded in financial knowledge.
@@ 2.2 Financial Reasoning
Real-world financial problems require precise numerical reasoning. FinQA (Chen et al., 2021) and ConvFinQA (Chen et al., 2022) were developed before explicit Chain-of-Thought reasoning became a standard evaluation target. As a result, they supervise arithmetic program generation but offer only weak step-level signals, yielding traces that are neither explicit nor verifiable. FinTextQA (Chen et al., 2024a) introduces long-form financial questions from textbooks and regulatory sources and focuses on explanatory retrieval rather than traceable computation. Bridging text and numerical reasoning, TAT-QA (Zhu et al., 2021) and MultiHiertt (Zhao et al., 2022) combine textual and tabular evidence, while DocMath-Eval (Zhao et al., 2023c) and FinanceMath (Zhao et al., 2023b) move toward interpretable symbolic evaluation. However, these datasets remain largely domain-agnostic and lack explicit, step-level supervision grounded in financial formulas. More recently, FinanceReasoning (Tang et al., 2025) improves answer-level numerical reliability by introducing executable Python solutions, but does not provide systematic verification of step-level reasoning alignment.
FinChain addresses this gap by introducing a symbolic, executable benchmark with explicit intermediate supervision and automatic alignment-based evaluation, spanning 58 topics across 12 financial domains.
@@ 3.1 Data Creation Process
We begin by identifying and defining financial domains based on established literature (Bodie et al., 2025) and expert input within the team, resulting in 12 distinct domains. Within each domain, we propose candidate financial topics with LLM assistance and curate them with financial experts, yielding a total of 58 topics (mean 4.8 per domain). The resulting taxonomy is illustrated in Figure 2. Following Mirzadeh et al. (2024), we instantiate each topic through parameterized symbolic templates that define both the question structure and an executable Chain-of-Thought solution grounded in domain-specific formulas. We implement these templates as executable Python programs that generate both intermediate reasoning steps and final answers, thereby enabling fully machine-verifiable evaluation. For each topic, we design five templates spanning three difficulty levels (two basic, two intermediate, and one advanced), where we control difficulty by the number and the complexity of required reasoning steps. Table 1 summarizes the dataset statistics and shows increasing reasoning depth across difficulty levels. An example symbolic template is shown in Figure 1. We generate the templates with LLM assistance and subsequently curate them with domain experts to ensure correctness, consistency, and balanced difficulty. We provide detailed prompt designs and generation procedures in Appendix A.1.
Statistic Basic Intermediate Advanced
#Templates 116 116 58
Avg. steps 2.01 2.97 3.90
Table 1: Dataset statistics of FinChain.
3.2 Data Validation and Expert Review
To ensure data quality and consistency, we apply a set of validation constraints covering numerical precision, unit consistency, input completeness, and reasoning step informativeness. Templates that fail validation are revised prior to expert review. A detailed description of the validation criteria is provided in Appendix A.3. We ask financial experts to review all validated templates following a calibrated annotation protocol. Specifically, all reviewers first participated in a pilot calibration phase, during which they jointly reviewed a shared subset of templates and discussed discrepancies to align on annotation standards. After this calibration phase, reviewers independently assessed the remaining templates, evaluating both the correctness of reasoning steps and the final numerical results under the agreed-upon criteria. Further details on annotator backgrounds, annotation procedures, and quality control are provided in Appendix A.4.
@@ 4 ChainEval
We propose ChainEval, an evaluation framework that jointly assesses reasoning-step alignment and final-answer correctness. Building on prior work on reasoning consistency (Lyu et al., 2023; Golovneva et al., 2023), our approach explicitly verifies intermediate results via step–answer matching while also checking the final numerical outcome.
@@ 4.1 Preliminaries
where $s^{*}_{i}$ and $\hat{s}_{j}$ denote individual reasoning steps in the gold and in the predicted solutions, respectively. Each step $s_{i}$ produces an intermediate result,
@@
representing the numerical or the symbolic value computed at that step.
To evaluate the reasoning faithfulness, we compare these sequences both semantically and numerically. In addition, we apply Dynamic Time Warping (DTW) to capture the global structural alignment between step sequences. DTW provides an order-preserving but flexible alignment that accommodates insertions, deletions, or small reordering of steps while maintaining the overall sequence coherence.
@@ 4.2 Reasoning Step Alignment
Each step is encoded using a sentence encoder $\mathrm{Enc}(\cdot)$, and the pairwise semantic similarity between the gold and the predicted steps is computed as
$\mathrm{SS}(s^{*}_{i},\hat{s}_{j})=\mathrm{cos}\!\left(\mathrm{Enc}(s^{*}_{i}),\mathrm{Enc}(\hat{s}_{j})\right),$ (3)
where $\mathrm{cos}(\cdot,\cdot)$ denotes the cosine similarity and $\mathrm{SS}\in[0,1]$.
@@
where $\mathbb{I}(\cdot)$ denotes the indicator function, and $\epsilon=0.05$ permits up to a 5% relative numerical deviation to account for rounding or error propagation. This design choice is motivated by financial auditing standards [^1] https://www.materialitytracker.net/standards/financial-thresholds/, in which materiality thresholds are commonly defined as between 5% and 10% of a base metric such as earnings, and deviations below 5% are generally considered immaterial.
@@
To capture global reasoning consistency, we align the two step sequences using Dynamic Time Warping (DTW). DTW searches for an optimal monotonic alignment path between (S,S^)($^{*},\hat{S}$) that minimizes the cumulative cost while respecting the step order. This formulation naturally handles local insertions, deletions, and step compressions, providing a structured measure of how closely the reasoning flows align.
DTWGate Alignment Score.
We transform the minimal DTW cost into a normalized similarity measure as follows:
$\mathrm{DTWGate}(S^{*},\hat{S})=1-\frac{\mathrm{Cost}_{\mathrm{DTW}}}{L_{\mathrm{path}}},$ (6)
where $\mathrm{Cost}_{\mathrm{DTW}}$ denotes the total alignment cost and $L_{\mathrm{path}}$ represents the length of the optimal alignment path. The resulting score lies in the range $[0,1]$, with higher values indicating stronger reasoning alignment between the gold and the predicted solutions.